Prove by induction, that the nth Fibonacci number can be found by the formula фт — фт Fr _ n V5 1-5 whereD
Answer #1

The Fibonacci sequence F_{n} is given by ,

F_{0} = F_{1}= 1 and F_{n+1} = F_{n} + F_{n-1}

Now we will use induction on n to prove that ,

В1 Ф — Ф Fn У5

Where , \Phi = \frac{1+\sqrt5}{2} , \phi = \frac{1-\sqrt5}{2}

Base step : For n =1

\frac{\Phi^{1} - \phi^{1}}{\sqrt5}

=\frac{\frac{1 +\sqrt5}{2}- \frac{1- \sqrt5}{2}}{\sqrt5}

=\frac{\sqrt5}{\sqrt5}

= 1

= F_{1}

So the formula is true for n =1 .

Induction hypothesis : Suppose the formupa is true for all n \leq m i.e., F_{m} = \frac{\Phi^m - \phi^{m}}{\sqrt5} , F_{m-1} = \frac{\Phi^{m-1}-\phi^{m-1}}{\sqrt5} .......(1)

Induction steps : For n = m+1

F_{m+1}

=F_{m} + F_{m-1}

=\frac{\Phi^m - \phi^{m}}{\sqrt5} +\frac{\Phi^{m-1}-\phi^{m-1}}{\sqrt5}

=\frac{1}{\sqrt5}[ \Phi^{m-1} (\Phi +1) - \phi^{m-1}(\phi +1)]

=\frac{1}{\sqrt5}[ \Phi^{m-1} (\frac{1+\sqrt5}{2}+1) -\phi^{m-1}(\frac{1-\sqrt5}{2}+1)]

=\frac{1}{\sqrt5}[ \Phi^{m-1} (\frac{3+\sqrt5}{2}) -\phi^{m-1}(\frac{3-\sqrt5}{2})]

:(Фт-1ф2 — фт-12? 5

=\frac{1}{\sqrt5}[ \Phi^{m+1} - \phi^{m+1}]

= F_{m+1}

So the formula is true for n=m+1 if we assume it is true for n \leq m also it id true for n=1 . Hence by induction on n the formula is true for all natural number n . Hence ,

В1 Ф — Ф Fn У5

.

.

.

.

If you any doubt or need more clarication at any step please comment.